Roscoe must play 27 commercials during his 1-hr radio show. Each commercial is either thirty seconds or sixty seconds long. If the total commercial time during that hour is 17 minutes, how many commercials of each type does Roscoe play?
Round answers to the nearest whole number.
30 second commercials_____
60 Second Commercials______

Answer by macston(5194)

.
T=number of 30 second commercials; S=number of 60 second commercials
.
T+S=27
T=27-S
.
17 min x 60 sec/min=1020 seconds
.
30T+60S=1020
30(27-S)+60S=1020
810-30S+60S=1020
30S=210
S=7
ANSWER 1: There were seven 60 second commercials.
.
T=27-S=27-7=20
ANSWER 2: There were twenty 30 second commercials.
.
CHECK:
30T+60S=1020
30(20)+60(7)=1020
600+420=1020
1020=1020
.